OpenStudy (anonymous):

what is the sum of the first eight terms of a geometric series whose first term is 3 and whose common ratio is 1/2

OpenStudy (campbell_st):

well you need the formula \[S_{8} = \frac{a(1 - (\frac{1}{2})^8}{1 - \frac{1}{2}}\]

OpenStudy (campbell_st):

all you need to do is substitute a = 3 and calculate an answer

OpenStudy (campbell_st):

opps should read \[S_{8} = \frac{a(1 -( \frac{1}{2})^8)}{1 -\frac{1}{2}}\]